# DisCoPatch: Taming Adversarially-driven Batch Statistics for Improved Out-of-Distribution Detection

### v4

@@ Abstract
Out-of-distribution (OOD) detection holds significant importance across many applications. While semantic and domain-shift OOD problems are well-studied, this work focuses on covariate shifts - subtle variations in the data distribution that can degrade machine learning performance. We hypothesize that detecting these subtle shifts can improve our understanding of in-distribution boundaries, ultimately improving OOD detection. In adversarial discriminators trained with Batch Normalization (BN), real and adversarial samples form distinct domains with unique batch statistics — a property we exploit for OOD detection. We introduce DisCoPatch, an unsupervised Adversarial Variational Autoencoder (VAE) framework that harnesses this mechanism. During inference, batches consist of patches from the same image, ensuring a consistent data distribution that allows the model to rely on batch statistics. DisCoPatch uses the VAE’s suboptimal outputs (generated and reconstructed) as negative samples to train the discriminator, thereby improving its ability to delineate the boundary between in-distribution samples and covariate shifts. By tightening this boundary, DisCoPatch achieves state-of-the-art results in public OOD detection benchmarks. The proposed model not only excels in detecting covariate shifts, achieving 95.5% AUROC on ImageNet-1K(-C), but also outperforms all prior methods on public Near-OOD (95.0%) benchmarks. With a compact model size of $\leq$ 25MB, it achieves high OOD detection performance at notably lower latency than existing methods, making it an efficient and practical solution for real-world OOD detection applications. The code is available at github.com/caetas/DisCoPatch.
@@ 3.3 Patching Strategy
In this setup, using the standard BatchNorm2D layer constrains batch processing during inference to a single image, because normalization must be applied only to features from the same distribution; mixing ID and OOD samples in a batch would distort the statistics and hinder effective separation. To address this, we introduce a custom layer, PatchNorm2D, which retains the weights and biases of BatchNorm2D but enables the normalization of multiple groups of patches simultaneously. Specifically, each batch consists of several groups of $N$ patches, where each group contains patches from the same image that are normalized together. This ensures that the model produces independent results per image while maintaining the benefits of batch normalization. The final anomaly score for an image is the mean of the scores of all patches within that same image. For the remainder of the paper, we refer to the model as DisCoPatch-$N$, indicating the number of patches per image used during inference. Both strategies are illustrated in Figure 4.
@@ 4.6 Batch Normalization Bias Analysis
MOODv2 [34] (BEiTv2) *85.0/58.1 *92.7/38.2 99.6/1.8 94.3/24.7 97.4/13.6 *70.5/73.9
SCALE [68] (ResNet-50) 77.4/67.7 85.4/51.8 98.0/9.5 97.6/11.9 94.0/28.2 *83.3/54.1
NNGuide [44] (RegNet) *84.7/54.7 *93.7/28.9 99.9/1.8 95.8/17.0 97.7/10.8 *78.5/61.6
NNGuide [44] (ResNet-50) *71.5/82.8 *80.5/69.8 96.9/14.3 90.4/27.4 92.4/35.4 *78.6/58.7
NNGuide [44] (MobileNetV2) *64.3/90.3 *75.5/80.2 80.2/70.2 88.3/40.6 82.9/63.3 *78.1/61.4
RankFeat [56] (ResNetv2-101) *89.4/47.9 *90.0/39.3 96.0/13.0 95.0/25.4 *92.4/33.3 *91.3/38.7
ASH [10] (ResNet-50) *74.1/80.6 *83.0/64.2 97.9/11.5 97.6/11.9 *92.8/32.7 *84.7/51.3
FDBD [37] (ResNet-50) *68.4/83.4 *81.4/66.1 97.6/12.4 98.0/10.7 *91.7/35.6 *82.2/57.8
@@ 5 Results
This section presents the key results of the conducted experiments. First, we demonstrate that BatchNorm exhibits a substantial reliance on batch statistics rather than the running mean and variance learned during training. Second, we showcase the results covering Near-OOD, Far-OOD, and Covariate Shift OOD detection performance. Additional detailed results are in the supplemental materials.
5.1 ImageNet-1K Benchmark
@@
The superiority of DisCoPatch in OOD detection extends beyond raw detection performance. As illustrated in Figure 5, DisCoPatch outperforms all other models in Near-OOD detection performance while reducing latency by as much as one order of magnitude. Even with only 16 patches per image, DisCoPatch outperforms all other models tested. Appendix F offers a detailed examination of the performance of additional patch counts. However, increasing the patch count beyond 64 did not result in any noticeable performance improvements. The performance gap between DisCoPatch and the second fastest model, NNGuide (MobileNetV2 backbone), is near 25%.
@@ 6 Discussion
Covariate Shift OOD. Validation on ImageNet-1K(-C) reveals that DisCoPatch achieves a substantial performance improvement over other models in this task. As hypothesized in Section 3, and supported by recent findings [36, 35], training a discriminator with VAE reconstructions enhances sensitivity to corruptions that diminish the high-frequency spectrum. This effect arises because reconstructed images generally lack high-frequency content, prompting the discriminator to classify this lack of content as "fake" and its presence as "real." In contrast, training the discriminator with generated images strengthens its ability to detect high-frequency amplification. DisCoPatch’s unique unsupervised training approach — exposing the discriminator to both generated and reconstructed patches — enables robust detection of both low- and high-frequency perturbations, leading to DisCoPatch’s consistent performance across diverse corruptions.
Near-OOD and Far-OOD. The covariate shift-focused training strategy effectively tightens the boundary between ID and OOD samples, improving detection performance on both Near-OOD and Far-OOD datasets. DisCoPatch outperforms existing models, achieving SOTA performance in Near-OOD detection. Although the same does not occur for Far-OOD, DisCoPatch’s performance is competitive, and the model is only beaten by far larger and slower models, like MOODv2 and NNGuide with the RegNet backbone.
Deployment. There are generally two main deployment scenarios for OOD detection algorithms: (1) The OOD detection algorithm is the primary focus, deployed as a standalone application. (2) The OOD detection algorithm operates alongside a main image processing algorithm, ensuring its safe and effective use. An OOD algorithm must be practical and effective in real-world scenarios, delivering strong detection performance while being highly deployable. Deployability should be assessed in the following aspects.
  1. 1.
    Accessibility: Evaluated by the compute requirements necessary for the algorithm.
  2. 2.
    Development Cycle: Measured by the time required for model training and deployment.
  3. 3.
    Inference Speed: The time it takes for the algorithm to make predictions during deployment.
  4. 4.
    Accuracy: The ability of the algorithm to provide highly accurate OOD detection.
An ideal OOD detection algorithm excels in all the above dimensions, ensuring it can be effectively utilized in various practical applications. As detailed in our results and Appendix C, DisCoPatch excels in all criteria. The model achieves SOTA OOD detection results, while utilizing substantially smaller and faster models. Table 4 indicates that DisCoPatch is 1 order of magnitude faster than the other evaluated models; being up to 12 times faster than MOODv2 and up to 19 times quicker than NNGuide. Training a tailored model is also efficient and flexible, as demonstrated in Table 3, making it a strong candidate for applications that require fast development cycles.
@@ 7 Limitations & Future Work
While current results demonstrate the effective use of batch statistics for OOD detection, a signal-processing-focused analysis of feature propagation and suppression at each layer could provide key insights. This could reveal patterns specific to different shifts, improving model interpretability. The planned extensions include the integration of several ResNet-based architectures that use BatchNorm2D layers in the discriminator. Beyond new discriminator architectures, future work should also assess more robust reconstructors (e.g., VQ-GAN) and generators (e.g., DDPMs).
@@ 8 Conclusion
This paper introduces DisCoPatch, an unsupervised, lightweight framework for OOD detection. DisCoPatch, unlike traditional VAE or GAN training objectives, uses a combination of reconstructed and generated images to address a wide range of frequency-spectrum perturbations. Furthermore, this study sheds light on the inherent bias in Batch Normalization toward batch statistics. DisCoPatch effectively exploits this bias for competitive OOD detection, particularly for Covariate Shift and Near-OOD detection.
DisCoPatch achieves SOTA performance on public OOD detection benchmarks. Specifically, it excels in Covariate Shift detection, achieving an AUROC of 95.5% on ImageNet-1K(-C) and outperforming all prior methods on Near-OOD detection with a score of 95.0%. DisCoPatch achieves these results with significantly lower latency (up to one order of magnitude) and a model size of $\leq$ 25MB, making it a viable option for real-time applications with limited resources.
@@ Appendix F Patch Count Imapct
MOODv2 88.9 97.1 70.5
SCALE 81.4 96.5 83.3
@@
DisCoPatch-4 (Proposed) 90.3+0.6 92.6-5.2 93.0+1.1
DisCoPatch-16 (Proposed) 94.3+4.6 96.0-1.8 96.5+3.6
DisCoPatch-32 (Proposed) 94.8+5.1 96.4-1.4 96.9+4.0
DisCoPatch-64 (Proposed) 95.1+5.4 96.6-1.2 97.2+4.3
DisCoPatch-128 (Proposed) 95.2+5.5 96.7-1.1 97.3+4.4
DisCoPatch-512 (Proposed) 95.3+6.1 96.8-1.0 97.4+4.5

### v5

@@ Abstract
Out-of-distribution (OOD) detection holds significant importance across many applications. While semantic and domain-shift OOD problems are well-studied, this work focuses on covariate shifts - subtle variations in the data distribution that can degrade machine learning performance. We hypothesize that detecting these subtle shifts can improve our understanding of in-distribution boundaries, ultimately improving OOD detection. In adversarial discriminators trained with Batch Normalization (BN), real and adversarial samples form distinct domains with unique batch statistics — a property we exploit for OOD detection. We introduce DisCoPatch, an unsupervised Adversarial Variational Autoencoder (VAE) framework that harnesses this mechanism. During inference, batches consist of patches from the same image, ensuring a consistent data distribution that allows the model to rely on batch statistics. DisCoPatch uses the VAE’s suboptimal outputs (generated and reconstructed) as negative samples to train the discriminator, thereby improving its ability to delineate the boundary between in-distribution samples and covariate shifts. By tightening this boundary, DisCoPatch achieves state-of-the-art results in public OOD detection benchmarks. The proposed model not only excels in detecting covariate shifts, achieving 95.5% AUROC on ImageNet-1K(-C), but also outperforms all prior methods on public Near-OOD (95.0%) benchmarks. With a compact model size of 25MB, it achieves high OOD detection performance at notably lower latency than existing methods, making it an efficient and practical solution for real-world OOD detection applications. The code is available at github.com/caetas/DisCoPatch.
@@ 3.3 Patching Strategy
In this setup, using the standard BatchNorm2D layer constrains batch processing during inference to a single image, since normalization must be applied only to features from the same distribution; mixing ID and OOD samples in a batch would distort the statistics and hinder effective separation. Therefore, we introduce a custom layer, PatchNorm2D, which retains the weights and biases of BatchNorm2D but enables the simultaneous normalization of multiple groups of patches. Specifically, each batch consists of several groups of $N$ patches, where each group contains patches from the same image that are normalized together. This ensures that the model produces independent results per image while maintaining the benefits of batch normalization. The final anomaly score for an image is the mean of the scores of all patches within that same image. For the remainder of the paper, we refer to the model as DisCoPatch-$N$, indicating the number of patches per image used during inference. Both strategies are illustrated in Figure 4.
@@ 4.6 Batch Normalization Bias Analysis
MOODv2 [34] (BEiTv2) 85.0/58.1 92.7/38.2 99.6/1.8 94.3/24.7 97.4/13.6 70.5/73.9
SCALE [68] (ResNet-50) 77.4/67.7 85.4/51.8 98.0/9.5 97.6/11.9 94.0/28.2 83.3/54.1
NNGuide [44] (RegNet) 84.7/54.7 93.7/28.9 99.9/1.8 95.8/17.0 97.7/10.8 78.5/61.6
NNGuide [44] (ResNet-50) 71.5/82.8 80.5/69.8 96.9/14.3 90.4/27.4 92.4/35.4 78.6/58.7
NNGuide [44] (MobileNetV2) 64.3/90.3 75.5/80.2 80.2/70.2 88.3/40.6 82.9/63.3 78.1/61.4
RankFeat [56] (ResNetv2-101) 89.4/47.9 90.0/39.3 96.0/13.0 95.0/25.4 92.4/33.3 91.3/38.7
ASH [10] (ResNet-50) 74.1/80.6 83.0/64.2 97.9/11.5 97.6/11.9 92.8/32.7 84.7/51.3
FDBD [37] (ResNet-50) 68.4/83.4 81.4/66.1 97.6/12.4 98.0/10.7 91.7/35.6 82.2/57.8
@@ 5 Results
The superiority of DisCoPatch extends beyond raw detection performance. As illustrated in Figure 5, DisCoPatch outperforms all other models in Near-OOD detection performance while reducing latency by up to an order of magnitude. Even with only 16 patches per image, DisCoPatch outperforms all other models tested. Appendix F offers a detailed examination of the performance of additional patch counts. However, increasing the patch count beyond 64 did not result in any noticeable performance improvements. The performance gap between DisCoPatch and the second fastest model, NNGuide (MobileNetV2), is near 25%.
@@ 6 Discussion
Covariate Shift OOD. Validation on ImageNet-1K(-C) reveals that DisCoPatch achieves a substantial improvement over other models in this task. As hypothesized in Section 3, and supported by recent findings [36, 35], training a discriminator with VAE reconstructions enhances sensitivity to corruptions that diminish the high-frequency spectrum. This effect arises because reconstructed images generally lack high-frequency content, prompting the discriminator to classify this lack of content as "fake" and its presence as "real." In contrast, training the discriminator with generated images strengthens its ability to detect high-frequency amplification. DisCoPatch’s unique unsupervised training approach, which exposes the discriminator to generated and reconstructed patches, enables robust detection of low- and high-frequency perturbations, leading to consistent performance across various corruptions.
Near-OOD and Far-OOD. The covariate shift-focused training strategy effectively tightens the boundary between ID and OOD samples, improving detection performance on both Near-OOD and Far-OOD datasets. DisCoPatch outperforms existing models, achieving SOTA performance in Near-OOD detection. Although the same does not occur for Far-OOD, DisCoPatch’s performance is competitive, and the model is only beaten by far larger and slower models.
Deployment. There are generally two main deployment scenarios for OOD detection algorithms: (1) The OOD detection algorithm is the primary focus, deployed as a standalone application. (2) The OOD detection algorithm operates alongside a main image processing algorithm, ensuring its safe and effective use. An OOD algorithm must be practical and effective in real-world scenarios, delivering strong detection performance while being highly deployable. Deployability should be assessed in the following aspects: (a) Accessibility, evaluated by the compute requirements necessary for the algorithm; (b) Development Cycle, measured by the time required for model training and deployment; (c) Inference Speed, which reflects the time it takes for the algorithm to make predictions during deployment; and (d) Accuracy, denoting the ability of the algorithm to provide highly accurate OOD detection.
An ideal OOD detection algorithm excels in all these dimensions, ensuring that it can be used effectively in various practical applications. As detailed in our results and Appendix C, DisCoPatch excels in all criteria. The model achieves SOTA OOD detection results, while utilizing a substantially smaller and faster model. Table 4 indicates that DisCoPatch is 1 order of magnitude faster than the other evaluated models; being up to 12 times faster than MOODv2 and up to 19 times quicker than NNGuide. Training a tailored model is also efficient and flexible, as demonstrated in Table 3, making it a strong candidate for applications that require fast development cycles.
@@ 7 Limitations & Future Work
While current results demonstrate the effective use of batch statistics for OOD detection, a signal-processing-focused analysis of feature propagation and suppression at each layer could provide key insights, improving model interpretability. The planned extensions include the integration of several ResNet-based architectures that use BatchNorm2D layers. Beyond discriminator architectures, future work should also assess more robust reconstructors (e.g., VQ-GAN) and generators (e.g., DDPMs).
@@ 8 Conclusion
This paper introduces DisCoPatch, an unsupervised lightweight framework for OOD detection. DisCoPatch, unlike traditional VAE or GAN training objectives, uses a combination of reconstructed and generated images to address a wide range of frequency-spectrum perturbations. Furthermore, this study highlights the inherent bias in Batch Normalization toward batch statistics. DisCoPatch effectively exploits this bias for competitive OOD detection, particularly for Covariate Shift and Near-OOD detection. DisCoPatch sets a new state-of-the-art in public OOD detection benchmarks, with 95.5% AUROC on Covariate Shift detection and 95.0% on Near-OOD detection, while maintaining low latency and a compact 25MB model, which makes it well suited for real-time, resource-constrained applications.
@@ Appendix F Patch Count Imapct
MOODv2 (BEiTv2) 88.9 97.1 70.5
SCALE (ResNet-50) 81.4 96.5 83.3
@@
DisCoPatch-4 (Proposed) +0.090.3+0.6 -0.092.6-5.2 +0.093.0+1.1
DisCoPatch-16 (Proposed) +0.094.3+4.6 -0.096.0-1.8 +0.096.5+3.6
DisCoPatch-32 (Proposed) +0.094.8+5.1 -0.096.4-1.4 +0.096.9+4.0
DisCoPatch-64 (Proposed) +0.095.1+5.4 -0.096.6-1.2 +0.097.2+4.3
DisCoPatch-128 (Proposed) +0.095.2+5.5 -0.096.7-1.1 +0.097.3+4.4
DisCoPatch-512 (Proposed) +0.095.3+5.6 -0.096.8-1.0 +0.097.4+4.5